40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 842     Entry time: Fri Aug 15 17:38:41 2008
Author: Yoichi 
Type: Update 
Category: SUS 
Subject: OSEM free swinging spectra before the pump down 
I ran an overnight measurement of the free swinging OSEM spectra.
The attm1 shows the results. Everything look ok except for the ITMY UL OSEM.
The time series from that OSEM was very noisy and had many spikes.
We suspected the cable from the satellite box to the computer rack because we disconnected the cable
when we tested a spare cable which was used to connect the spare OSEMs to the PRM suspension in the clean room.
Janne remembered when she put the cable back, she trusted the latch on the connector and did not push it in too hard.
However, Rob suggested the latch does not work well. So she pushed the connector again. Then the signal from
the ITMY UL OSEM got back to normal.
The second attachment shows the ITMY spectra after the cable push.
We decided to pump down after confirming this.

There are still a lot of extra peaks especially in the suspensions in the BS chamber.
These may be inter modulations (by the non-linearities of the OSEMs) of the modes of the multiple
suspensions sitting on the same stack.
Attachment 1: 2008-8-15.pdf  1.546 MB  | Hide | Hide all | Show all
2008-8-15.pdf 2008-8-15.pdf 2008-8-15.pdf 2008-8-15.pdf 2008-8-15.pdf 2008-8-15.pdf 2008-8-15.pdf 2008-8-15.pdf
Attachment 2: ITMY2.pdf  156 kB  | Show | Hide all | Show all
ELOG V3.1.3-